OpenStudy (katelyn1273):

I need help on a few questions that I am stuck on I will be putting down what I think it is but I am not 100% sure. #1 What method is the most efficient to use to solve X^2+4x+3=0 A: Factoring B: Isolating the X^2 term and finding the square root of both sides. C: Using the Quadratic Formula D: All three methods would be efficient. I say C #2 What method is the most efficient to use to solve X^2+5x-6=0 A: Factoring B: Isolating the X^2 term and finding the square root of both sides. C: Using the quadratic formula D: All three methods would be efficient I say D

OpenStudy (katelyn1273):

Here are the other 2 questions. #3 Which method is the most efficient to use to solve X^2-7=0 A: factoring B:Isolating the X^2 term and finding the square root of both sides. C: Using the quadratic formula. D: All three methods would be efficient. I say A #4 Lisa is planning to build a rectangular deck that is 10 feet longer than it is wide. Which equation would Lisa use to determine the width of the deck when the area is 375 square feet. A: 10w^2=375 B: w^2+10w=375 C: W+10w=375 D: W^2-10w=375 im not sure which one it is sorry.